Snoop bought a new house, the price of the house is $5,000,000. Snoop will put down 20% using a 30-year mortgage with an APR of 3.375%? What is the balance on Snoop's mortgage after 5 years?
Answer #1

Loan amount = (1-20%)*5,000,000 = $4000000

We first find the annual interest payment

Using a financial calculator

PV = -4000000

FV = 0

N = 30

I/Y = 3.375

cpt PMT, we get PMT = 214092.64

Now, we find the loan amount remaining after 5 years

Using a financial calculator

PV = -4000000

PMT = 214092.64

N = 5

I/Y = 3.375

cpt FV, we get FV = 3576926.83

Hence, balance on Snoop's mortgage after 5 years = $3576926.83